# Square Functions Controlling Smoothness and Higher-Order Quantitative Rectifiability

### v1

@@ 3. $I_{\alpha}(BMO)$ Implies Square Function Bounds
$VI=\int_{|x-z|<100R}\left(\int_{|x-y|\geq 10R}\frac{(\nabla f(z)-\nabla f(x))\cdot(x-y)}{|x-y|^{d+\alpha}}\phi\left(\frac{y-z}{R}\right)\,\mathrm{d}y\right)^{2}\,\mathrm{d}x\\ \lesssim_{[\nabla f]_{\alpha-1}^{2}}\int_{|x-z|<100R}\left(\int_{|x-y|\geq 10R}\frac{R^{\alpha-1}}{|x-y|^{d+\alpha-1}}\phi\left(\frac{y-z}{R}\right)\,\mathrm{d}y\right)^{2}\,\mathrm{d}x\\ \approx_{[\nabla f]_{\alpha-1}^{2}}R^{2\alpha-2}\int_{|x-z|<100R}\left(\int_{CR}^{\infty}r^{-\alpha}\,\mathrm{d}r\right)^{2}\,\mathrm{d}x\lesssim R^{d}$
@@
$\frac{(\nabla f(z)-\nabla f(x))\cdot(x-y)}{|x-y|^{d+\alpha}}1_{{}_{|x-y|\leq 10R}}$
@@
$\int_{|x-z|<100R}\left(p.v.\int_{|x-y|\leq 10R}\frac{(\nabla f(z)-\nabla f(x))\cdot(x-y)}{|x-y|^{d+\alpha}}\left(\phi\left(\frac{y-z}{R}\right)-\phi\left(\frac{x-z}{R}\right)\right)\,\mathrm{d}y\right)^{2}\,\mathrm{d}x$
@@
$K(x,y)=\frac{f(y)-f(x)-\nabla f(x)\cdot(y-x)}{|x-y|^{d+\alpha}}$
We claim that $K(x,y)$ satisfies the standard Calderón-Zygmund kernel estimates. The size estimate is easy to verify. For fixed $x,y$, $x\neq y$, and $h$ such that $2|h|\leq|x-y|$, we have
$|K(x,y+h)-K(x,y)|\leq\left|\frac{\big{(}f(y+h)-f(x)-\nabla f(x)\cdot(y+h-x)\big{)}}{|x-y-h|^{d+\alpha}}-\frac{\big{(}f(y)-f(x)-\nabla f(x)\cdot(y-x)\big{)}}{|x-y-h|^{d+\alpha}}\right|\\ +\left|\left(\frac{1}{|x-y-h|^{d+\alpha}}-\frac{1}{|x-y|^{d+\alpha}}\right)\big{(}f(y)-f(x)-\nabla f(x)\cdot(y-x)\big{)}\right|\\ \leq\frac{|f(y+h)-f(y)-\nabla f(y)\cdot h|}{|x-y-h|^{d+\alpha}}+\frac{|(\nabla f(y)-\nabla f(x))\cdot h|}{|x-y-h|^{d+\alpha}}\\ +\left|\left(\frac{1}{|x-y-h|^{d+\alpha}}-\frac{1}{|x-y|^{d+\alpha}}\right)\big{(}f(y)-f(x)-\nabla f(x)\cdot(y-x\big{)}\right|$
@@
$K^{*}(x,y)=-\frac{f(x)-f(y)-\nabla f(y)\cdot(x-y)}{|x-y|^{d+\alpha}}$
@@
$T[1](x)=p.v.\int_{\mathbb{R}^{d}}K(x,y)\,\mathrm{d}y\in BMO$
For to see that $T^{*}(1)\in BMO$, notice that
$p.v.\int_{\mathbb{R}^{d}}\frac{f(x)-f(y)}{|x-y|^{d+\alpha}}\,\mathrm{d}y\in BMO$
So, we just need to check that
$p.v.\int\frac{x-y}{|x-y|^{d+\alpha}}\nabla f(y)\,\mathrm{d}y\in BMO$
But it is clear that the above expression is a constant times $D_{\alpha}f$, at least if one first checks the distributional identity
$\bigg{(}p.v.\frac{x_{j}-y_{j}}{|x-y|^{d+\alpha}}\hat{\bigg{)}}(\xi)=iC(d,\alpha)\xi_{j}|\xi|^{\alpha-2}$
@@
$|\langle\varphi_{1},T^{*}\varphi_{2}\rangle|\leq Cr^{n}(\|\varphi_{1}\|_{\infty}+r\|\nabla\varphi_{1}\|_{\infty})(\|\varphi_{1}\|_{\infty}+r\|\nabla\varphi_{2}\|_{\infty})$
for some constant $C>0$. This was essentially proved in Lemma 4.3 of [H]. Here the author proves that the same operator, but with $\alpha=1$, satisfies the weak boundedness property if $\nabla f\in BMO$. In our setting, the exact same argument works, the only change required being that instead of obtaining bounds in terms of $\|\nabla f\|_{*}$, we obtain bounds in terms of $[\nabla f]_{\alpha-1}$ by using Campanato space estimates. In particular, $T^{*}$ satisfies the weak boundedness property above with $C\approx[\nabla f]_{\alpha-1}$. So, by the $T(1)$ theorem, $T^{*}:L^{2}(\mathbb{R}^{d})\rightarrow L^{2}(\mathbb{R}^{d})$ with $\|T^{*}\|_{L^{2}\rightarrow L^{2}}\leq C([\nabla f]_{\alpha-1}+\|D_{\alpha}f\|_{*})$. This mean that $T$ is $L^{2}\rightarrow L^{2}$ bounded with the same constant, so
$VIII\leq C([\nabla f]_{\alpha-1}^{2}+\|D_{\alpha}f\|_{*}^{2})R^{d}$
We conclude the proof of the theorem by noting that in [Str] the author proves $[\nabla f]_{\alpha-1}\lesssim\|D_{\alpha}f\|_{*}$. ∎
@@
$g(x):=\phi\left(\frac{x-z}{R}\right)\big{(}f(x)-f(z)\big{)}$
@@
(3.5) $\int_{\mathbb{R}^{d}}\big{|}|\xi|^{\alpha}\hat{g}(\xi)\big{|}^{2}\,\mathrm{d}\xi\leq CR^{d}$
@@
The proof of this Proposition is extremely similar to the proof of Proposition 3.1. It only differs in minor ways, and these differences actually make the proof even easier. We provide an outline for the reader. We remark that, whenever bounds were obtained for terms on the order of $[\nabla f]_{\alpha-1}^{2}$ in the proof of the last Proposition, the corresponding bounds in the proof of this proposition will be on the order of $[\nabla f]_{\alpha}^{2}.$
@@
$g(x)=\phi\left(\frac{x-z}{R}\right)(f(x)-f(z))$
@@
$\int_{\mathbb{R}^{d}}\big{|}|\xi|^{\alpha}\hat{g}(\xi)\big{|}^{2}\,\mathrm{d}\xi\approx\int_{\mathbb{R}^{d}}\left(p.v.\int_{\mathbb{R}^{d}}\frac{g(x)-g(y)}{|x-y|^{d+\alpha}}\,\mathrm{d}y\right)^{2}\,\mathrm{d}x$
@@
$I=\int_{|x-z|\geq 100R}\left(p.v.\int_{\mathbb{R}^{d}}\frac{g(x)-g(y)}{|x-y|^{d+\alpha}}\,\mathrm{d}y\right)^{2}\,\mathrm{d}x$
$II=\int_{|x-z|\leq 100R}\left(p.v.\int_{\mathbb{R}^{d}}\frac{g(x)-g(y)}{|x-y|^{d+\alpha}}\,\mathrm{d}y\right)^{2}\,\mathrm{d}x$
@@
$\int_{|x-z|\geq 100R}\left(p.v.\int_{\mathbb{R}^{d}}\frac{\phi\left(\frac{y-z}{R}\right)\big{(}f(y)-f(z)\big{)}}{|x-y|^{d+\alpha}}\,\mathrm{d}y\right)^{2}\,\mathrm{d}x$
@@
$\int_{|x-z|<100R}\left(p.v.\int_{\mathbb{R}^{d}}\frac{f(x)-f(y)}{|x-y|^{d+\alpha}}\phi\left(\frac{y-z}{R}\right)\,\mathrm{d}y\right)^{2}\,\mathrm{d}x$
@@
$\left|\phi\left(\frac{x-z}{R}\right)-\phi\left(\frac{y-z}{R}\right)\right|\lesssim\frac{|x-y|}{R}$
@@
(3.6) $\overline{\nu}^{f}_{1}(x,r)$ $:=\left(\fint_{B_{r}(x)}\left(\frac{|f(y)-(f\ast\psi_{r})(x)-\nabla(f\ast\psi_{r})(x)\cdot(y-x)|}{r}\right)^{2}\,\mathrm{d}y\right)^{\frac{1}{2}}$
(3.7) $\overline{\nu}^{f}_{0}(x,r)$ $:=\left(\fint_{B_{r}(x)}\left(\frac{|f(y)-(f\ast\psi_{r})(x)|}{r}\right)^{2}\,\mathrm{d}y\right)^{\frac{1}{2}}$
@@
$\int_{0}^{\infty}\int_{\mathbb{R}^{d}}\left(\frac{\overline{\nu}_{\lfloor\alpha\rfloor}^{f}(x,r)}{r^{\alpha-1}}\right)^{2}\frac{\,\mathrm{d}x\,\mathrm{d}r}{r}\lesssim\int_{\mathbb{R}^{d}}|\hat{f}(\xi)|\xi|^{\alpha}|^{2}\,\mathrm{d}\xi$
@@
(3.9) $\int_{0}^{\infty}\int_{\mathbb{R}^{d}}\left(\frac{\overline{\nu}_{0}^{f}(x,r)}{r^{\alpha-1}}\right)^{2}\frac{\,\mathrm{d}x\,\mathrm{d}r}{r}\\ =\int_{0}^{\infty}\int_{\mathbb{R}^{d}}\int_{B_{r}(x)}|f(y)-f\ast\psi_{r}(x)|^{2}\,\mathrm{d}y\,\mathrm{d}x\frac{\,\mathrm{d}r}{r^{d+1+2\alpha}}\\ =\int_{0}^{\infty}\int_{B_{r}(0)}\int_{\mathbb{R}^{d}}|f(x-y)-f\ast\psi_{r}(x)|^{2}\,\mathrm{d}x\,\mathrm{d}y\frac{\,\mathrm{d}r}{r^{d+1+2\alpha}}\\ =\int_{0}^{\infty}\int_{B_{r}(0)}\int_{\mathbb{R}^{d}}|\hat{f}(\xi)e^{-2\pi iy\cdot\xi}-\hat{f}(\xi)\hat{\psi}(r\xi)|^{2}\,\mathrm{d}\xi\,\mathrm{d}y\frac{\,\mathrm{d}r}{r^{d+1+2\alpha}}\\ =\int_{\mathbb{R}^{d}}|\hat{f}(\xi)|^{2}\int_{0}^{\infty}\int_{B_{r}(0)}|e^{-2\pi iy\cdot\xi}-\hat{\psi}(r\xi)|^{2}\,\mathrm{d}y\frac{\,\mathrm{d}r}{r^{d+1+2\alpha}}\,\mathrm{d}\xi$
@@
$\int_{\mathbb{R}^{d}}|\hat{f}(\xi)|^{2}|\xi|^{d+2\alpha}\int_{0}^{\infty}\int_{B_{\frac{r}{|\xi|}(0)}}|e^{-2\pi iy\cdot\xi}-\hat{\psi}(r)|^{2}\,\mathrm{d}y\frac{\,\mathrm{d}r}{r^{d+2\alpha}}\,\mathrm{d}\xi\\ \approx\int_{\mathbb{R}^{d}}|\hat{f}(\xi)|\xi|^{\alpha}|^{2}\int_{0}^{\infty}\int_{B_{r}(0)}|e^{-2\pi iy\cdot\frac{\xi}{|\xi|}}-\hat{\psi}(r)|^{2}\,\mathrm{d}y\frac{\,\mathrm{d}r}{r^{d+1+2\alpha}}\,\mathrm{d}\xi$
@@
We are going to show that the inner integrand $|e^{-2\pi iy\cdot\frac{\xi}{|\xi|}}-\hat{\psi}(r)|$ is bounded on the order of $\min\{r,1\}$. First we show that the integrand is less than $cr$ for $0\leq r\leq 1$. We use the triangle inequality
@@
$\int_{0}^{\infty}\int_{\mathbb{R}^{d}}\left(\frac{\nu_{0}^{f}(x,r)}{r^{\alpha-1}}\right)^{2}\frac{\,\mathrm{d}x\,\mathrm{d}r}{r}\lesssim\int_{\mathbb{R}^{d}}|\hat{f}(\xi)|\xi|^{\alpha}|^{2}\,\mathrm{d}\xi$
We now prove the Proposition in the case that $21\leq\alpha<$ 2. Taking almost identical steps to those taken in 3.9, we find
@@
$\leq\int_{\mathbb{R}^{d}}|\hat{f}(\xi)|\xi|^{\alpha}|^{2}\int_{0}^{\infty}\int_{B_{r}(0)}|e^{-2\pi iy\cdot\frac{\xi}{|\xi|}}-2\pi iy\cdot\frac{\xi}{|\xi|}\hat{\psi}(r)-\hat{\psi}(r)|^{2}\,\mathrm{d}y\frac{\,\mathrm{d}r}{r^{d+1+2\alpha}}\,\mathrm{d}\xi$
@@
(3.10) $|e^{-2\pi iy\cdot\frac{\xi}{|\xi|}}-2\pi iy\cdot\frac{\xi}{|\xi|}\hat{\psi}(r)-\hat{\psi}(r)|\leq|e^{-2\pi iy\cdot\frac{\xi}{|\xi|}}(1-\hat{\psi}(r))|+|\hat{\psi}(r)||e^{-2\pi iy\cdot\frac{\xi}{|\xi|}}-2\pi iy\cdot\frac{\xi}{|\xi|}-1|$
@@
$|\hat{\psi}(r)||e^{-2\pi iy\cdot\frac{\xi}{|\xi|}}-2\pi iy\cdot\frac{\xi}{|\xi|}-1|\lesssim|\hat{\psi}(r)|(1+r)\lesssim\frac{1+r}{1+r^{2}}\leq 1$
@@
$|\hat{\psi}(r)||e^{-2\pi iy\cdot\frac{\xi}{|\xi|}}-2\pi iy\cdot\frac{\xi}{|\xi|}-1|\lesssim|y|^{2}\leq r^{2}$
@@
$\lesssim\int_{0}^{1}r^{d+4}r^{-d-1-2\alpha}\,\mathrm{d}r+\int_{1}^{\infty}\frac{\,\mathrm{d}r}{r^{1+2\alpha}}<\infty$
@@
(3.12) $\int_{0}^{R}\int_{B_{R}(z)}\left(\frac{\overline{\nu}_{\lfloor\alpha\rfloor}^{f}(x,r)}{r^{\alpha-1}}\right)^{2}\frac{\,\mathrm{d}x\,\mathrm{d}r}{r}\leq CR^{d}$
@@
$g(x):=\phi\bigg{(}\frac{x-z}{R}\bigg{)}\big{(}f(x)-f(z)+\nabla(f\ast\varphi_{R})(z)\cdot(x-z)\big{)}$
@@
$\int_{0}^{R}\int_{B_{R}(z)}\left(\overline{\nu}_{1}^{f}(x,r)\right)^{2}\frac{\,\mathrm{d}x\,\mathrm{d}r}{r}=\int_{0}^{R}\int_{B_{R}(z)}\left(\overline{\nu}_{1}^{g}(x,r)\right)^{2}\frac{\,\mathrm{d}x\,\mathrm{d}r}{r}$

### v2

@@ 3. $I_{\alpha}(BMO)$ Implies Square Function Bounds
$VI=\int_{|x-z|<100R}\left(\int_{|x-y|\geq 10R}\frac{(\nabla f(z)-\nabla f(x))\cdot(x-y)}{|x-y|^{d+\alpha}}\phi\left(\frac{y-z}{R}\right)\,\mathrm{d}y\right)^{2}\,\mathrm{d}x\\ \lesssim_{[\nabla f]_{\alpha-1}^{2}}\int_{|x-z|<100R}\left(\int_{|x-y|\geq 10R}\frac{R^{\alpha-1}}{|x-y|^{d+\alpha-1}}\phi\left(\frac{y-z}{R}\right)\,\mathrm{d}y\right)^{2}\,\mathrm{d}x\\ \approx_{[\nabla f]_{\alpha-1}^{2}}R^{2\alpha-2}\int_{|x-z|<100R}\left(\int_{CR}^{\infty}r^{-\alpha}\,\mathrm{d}r\right)^{2}\,\mathrm{d}x\lesssim R^{d}.$
@@
$\frac{(\nabla f(z)-\nabla f(x))\cdot(x-y)}{|x-y|^{d+\alpha}}1_{{}_{|x-y|\leq 10R}}.$
@@
$\int_{|x-z|<100R}\left(p.v.\int_{|x-y|\leq 10R}\frac{(\nabla f(z)-\nabla f(x))\cdot(x-y)}{|x-y|^{d+\alpha}}\left(\phi\left(\frac{y-z}{R}\right)-\phi\left(\frac{x-z}{R}\right)\right)\,\mathrm{d}y\right)^{2}\,\mathrm{d}x.$
@@
$K(x,y)=\frac{f(y)-f(x)-\nabla f(x)\cdot(y-x)}{|x-y|^{d+\alpha}}.$
We claim that $K(x,y)$ satisfies the standard Calderón-Zygmund kernel estimates. We will verify
(size) $|K(x,y)|$ $\lesssim\frac{1}{|x-y|^{d}},\,\,\text{$x\neq y$},$
[…]
We will need both of these estimates to apply the $T(1)$ Theorem below. For fixed $x,y$, $x\neq y$, and $h$ such that $2|h|\leq|x-y|$, we have
$|K(x,y+h)-K(x,y)|\leq\left|\frac{\big{(}f(y+h)-f(x)-\nabla f(x)\cdot(y+h-x)\big{)}}{|x-y-h|^{d+\alpha}}-\frac{\big{(}f(y)-f(x)-\nabla f(x)\cdot(y-x)\big{)}}{|x-y-h|^{d+\alpha}}\right|\\ +\left|\left(\frac{1}{|x-y-h|^{d+\alpha}}-\frac{1}{|x-y|^{d+\alpha}}\right)\big{(}f(y)-f(x)-\nabla f(x)\cdot(y-x)\big{)}\right|\\ \leq\frac{|f(y+h)-f(y)-\nabla f(y)\cdot h|}{|x-y-h|^{d+\alpha}}+\frac{|(\nabla f(y)-\nabla f(x))\cdot h|}{|x-y-h|^{d+\alpha}}\\ +\left|\left(\frac{1}{|x-y-h|^{d+\alpha}}-\frac{1}{|x-y|^{d+\alpha}}\right)\big{(}f(y)-f(x)-\nabla f(x)\cdot(y-x\big{)}\right|.$
@@
$K^{*}(x,y)=-\frac{f(x)-f(y)-\nabla f(y)\cdot(x-y)}{|x-y|^{d+\alpha}}.$
@@
$T[1](x)=p.v.\int_{\mathbb{R}^{d}}K(x,y)\,\mathrm{d}y\in BMO,\,\,\,\|T[1]\|_{*}=\|D_{\alpha}f\|_{*}.$
To see that $T^{*}[1]\in BMO$ with $\|T^{*}(1)\|_{*}=\|D_{\alpha}f\|_{*}$, notice that $T^{*}[1]=-T[1]$. So, we just need to check that
$p.v.\int\frac{x-y}{|x-y|^{d+\alpha}}\nabla f(y)\,\mathrm{d}y\in BMO,$
with an appropriate bound on the $BMO$ norms. But it is clear that the above expression is a constant times $D_{\alpha}f$, at least if one first checks the distributional identity
$\bigg{(}p.v.\frac{x_{j}-y_{j}}{|x-y|^{d+\alpha}}\hat{\bigg{)}}(\xi)=iC(d,\alpha)\xi_{j}|\xi|^{\alpha-2}.$
@@
(3.4) $|\langle\varphi_{1},T^{*}\varphi_{2}\rangle|\leq Cr^{n}(\|\varphi_{1}\|_{\infty}+r\|\nabla\varphi_{1}\|_{\infty})(\|\varphi_{1}\|_{\infty}+r\|\nabla\varphi_{2}\|_{\infty})$
for some constant $C>0$. This was essentially proved in Lemma 4.3 of [H]. Here the author proves that the same operator, but with $\alpha=1$, satisfies the weak boundedness property if $\nabla f\in BMO$. To be precise, the author shows that inequality (3.4) holds with $C\approx\|\nabla f\|_{*}$. In our setting, the exact same argument works, the only change required being that instead of obtaining bounds in terms of $\|\nabla f\|_{*}$, we obtain bounds in terms of $[\nabla f]_{\alpha-1}$ by using Campanato space estimates. In particular, $T^{*}$ satisfies the weak boundedness property above with $C\approx[\nabla f]_{\alpha-1}$. So, by the $T(1)$ Theorem, $T^{*}:L^{2}(\mathbb{R}^{d})\rightarrow L^{2}(\mathbb{R}^{d})$ with $\|T^{*}\|_{L^{2}\rightarrow L^{2}}\leq C([\nabla f]_{\alpha-1}+\|D_{\alpha}f\|_{*})$. This mean that $T$ is $L^{2}\rightarrow L^{2}$ bounded with the same constant, so
$VIII\leq C([\nabla f]_{\alpha-1}^{2}+\|D_{\alpha}f\|_{*}^{2})R^{d}.$
We conclude the proof by recalling that $[\nabla f]_{\alpha-1}\lesssim\|D_{\alpha}f\|_{*}$. ∎
@@
$g(x):=\phi\left(\frac{x-z}{R}\right)\big{(}f(x)-f(z)\big{)}.$
@@
(3.6) $\int_{\mathbb{R}^{d}}\big{|}|\xi|^{\alpha}\hat{g}(\xi)\big{|}^{2}\,\mathrm{d}\xi\leq CR^{d},$
@@
The proof of this Proposition is extremely similar to the proof of Proposition 3.1. It only differs in minor ways, and these differences actually make the proof even easier. We provide an outline for the reader. We remark that, whenever bounds were obtained for terms on the order of $[\nabla f]_{\alpha-1}^{2}$ in the proof of the last Proposition, the corresponding bounds in the proof of this proposition will be on the order of $[f]_{\alpha}^{2}.$
@@
$g(x)=\phi\left(\frac{x-z}{R}\right)(f(x)-f(z)).$
@@
$\int_{\mathbb{R}^{d}}\big{|}|\xi|^{\alpha}\hat{g}(\xi)\big{|}^{2}\,\mathrm{d}\xi\approx\int_{\mathbb{R}^{d}}\left(p.v.\int_{\mathbb{R}^{d}}\frac{g(x)-g(y)}{|x-y|^{d+\alpha}}\,\mathrm{d}y\right)^{2}\,\mathrm{d}x.$
@@
$I=\int_{|x-z|\geq 100R}\left(p.v.\int_{\mathbb{R}^{d}}\frac{g(x)-g(y)}{|x-y|^{d+\alpha}}\,\mathrm{d}y\right)^{2}\,\mathrm{d}x,$
$II=\int_{|x-z|\leq 100R}\left(p.v.\int_{\mathbb{R}^{d}}\frac{g(x)-g(y)}{|x-y|^{d+\alpha}}\,\mathrm{d}y\right)^{2}\,\mathrm{d}x.$
@@
$\int_{|x-z|\geq 100R}\left(p.v.\int_{\mathbb{R}^{d}}\frac{\phi\left(\frac{y-z}{R}\right)\big{(}f(y)-f(z)\big{)}}{|x-y|^{d+\alpha}}\,\mathrm{d}y\right)^{2}\,\mathrm{d}x.$
@@
$\int_{|x-z|<100R}\left(p.v.\int_{\mathbb{R}^{d}}\frac{f(x)-f(y)}{|x-y|^{d+\alpha}}\phi\left(\frac{y-z}{R}\right)\,\mathrm{d}y\right)^{2}\,\mathrm{d}x.$
@@
$\left|\phi\left(\frac{x-z}{R}\right)-\phi\left(\frac{y-z}{R}\right)\right|\lesssim\frac{|x-y|}{R}.$
@@
(3.7) $\overline{\nu}^{f}_{1}(x,r)$ $:=\left(\fint_{B_{r}(x)}\left(\frac{|f(y)-(f\ast\psi_{r})(x)-\nabla(f\ast\psi_{r})(x)\cdot(y-x)|}{r}\right)^{2}\,\mathrm{d}y\right)^{\frac{1}{2}},$
(3.8) $\overline{\nu}^{f}_{0}(x,r)$ $:=\left(\fint_{B_{r}(x)}\left(\frac{|f(y)-(f\ast\psi_{r})(x)|}{r}\right)^{2}\,\mathrm{d}y\right)^{\frac{1}{2}}.$
@@
$\int_{0}^{\infty}\int_{\mathbb{R}^{d}}\left(\frac{\overline{\nu}_{\lfloor\alpha\rfloor}^{f}(x,r)}{r^{\alpha-1}}\right)^{2}\frac{\,\mathrm{d}x\,\mathrm{d}r}{r}\lesssim\int_{\mathbb{R}^{d}}|\hat{f}(\xi)|\xi|^{\alpha}|^{2}\,\mathrm{d}\xi.$
@@
(3.10) $\int_{0}^{\infty}\int_{\mathbb{R}^{d}}\left(\frac{\overline{\nu}_{0}^{f}(x,r)}{r^{\alpha-1}}\right)^{2}\frac{\,\mathrm{d}x\,\mathrm{d}r}{r}\\ =\int_{0}^{\infty}\int_{\mathbb{R}^{d}}\int_{B_{r}(x)}|f(y)-f\ast\psi_{r}(x)|^{2}\,\mathrm{d}y\,\mathrm{d}x\frac{\,\mathrm{d}r}{r^{d+1+2\alpha}}\\ =\int_{0}^{\infty}\int_{B_{r}(0)}\int_{\mathbb{R}^{d}}|f(x-y)-f\ast\psi_{r}(x)|^{2}\,\mathrm{d}x\,\mathrm{d}y\frac{\,\mathrm{d}r}{r^{d+1+2\alpha}}\\ =\int_{0}^{\infty}\int_{B_{r}(0)}\int_{\mathbb{R}^{d}}|\hat{f}(\xi)e^{-2\pi iy\cdot\xi}-\hat{f}(\xi)\hat{\psi}(r\xi)|^{2}\,\mathrm{d}\xi\,\mathrm{d}y\frac{\,\mathrm{d}r}{r^{d+1+2\alpha}}\\ =\int_{\mathbb{R}^{d}}|\hat{f}(\xi)|^{2}\int_{0}^{\infty}\int_{B_{r}(0)}|e^{-2\pi iy\cdot\xi}-\hat{\psi}(r\xi)|^{2}\,\mathrm{d}y\frac{\,\mathrm{d}r}{r^{d+1+2\alpha}}\,\mathrm{d}\xi.$
@@
$\int_{\mathbb{R}^{d}}|\hat{f}(\xi)|^{2}|\xi|^{d+2\alpha}\int_{0}^{\infty}\int_{B_{\frac{r}{|\xi|}(0)}}|e^{-2\pi iy\cdot\xi}-\hat{\psi}(r)|^{2}\,\mathrm{d}y\frac{\,\mathrm{d}r}{r^{d+2\alpha}}\,\mathrm{d}\xi\\ \approx\int_{\mathbb{R}^{d}}|\hat{f}(\xi)|\xi|^{\alpha}|^{2}\int_{0}^{\infty}\int_{B_{r}(0)}|e^{-2\pi iy\cdot\frac{\xi}{|\xi|}}-\hat{\psi}(r)|^{2}\,\mathrm{d}y\frac{\,\mathrm{d}r}{r^{d+1+2\alpha}}\,\mathrm{d}\xi.$
@@
We are going to show that the inner integrand $|e^{-2\pi iy\cdot\frac{\xi}{|\xi|}}-\hat{\psi}(r)|$ is bounded on the order of $\min\{r,1\}$ when $y\in B_{r}(0)$. First we show that the integrand is less than $cr$ for $0\leq r\leq 1$. We use the triangle inequality
@@
$\int_{0}^{\infty}\int_{\mathbb{R}^{d}}\left(\frac{\nu_{0}^{f}(x,r)}{r^{\alpha-1}}\right)^{2}\frac{\,\mathrm{d}x\,\mathrm{d}r}{r}\lesssim\int_{\mathbb{R}^{d}}|\hat{f}(\xi)|\xi|^{\alpha}|^{2}\,\mathrm{d}\xi.$
We now prove the Proposition in the case that $21\leq\alpha<$ 2. Taking almost identical steps to those taken in 3.10, we find
@@
$\leq\int_{\mathbb{R}^{d}}|\hat{f}(\xi)|\xi|^{\alpha}|^{2}\int_{0}^{\infty}\int_{B_{r}(0)}|e^{-2\pi iy\cdot\frac{\xi}{|\xi|}}-2\pi iy\cdot\frac{\xi}{|\xi|}\hat{\psi}(r)-\hat{\psi}(r)|^{2}\,\mathrm{d}y\frac{\,\mathrm{d}r}{r^{d+1+2\alpha}}\,\mathrm{d}\xi.$
@@
(3.11) $|e^{-2\pi iy\cdot\frac{\xi}{|\xi|}}-2\pi iy\cdot\frac{\xi}{|\xi|}\hat{\psi}(r)-\hat{\psi}(r)|\leq|e^{-2\pi iy\cdot\frac{\xi}{|\xi|}}(1-\hat{\psi}(r))|+|\hat{\psi}(r)||e^{-2\pi iy\cdot\frac{\xi}{|\xi|}}-2\pi iy\cdot\frac{\xi}{|\xi|}-1|.$
@@
$|\hat{\psi}(r)||e^{-2\pi iy\cdot\frac{\xi}{|\xi|}}-2\pi iy\cdot\frac{\xi}{|\xi|}-1|\lesssim|\hat{\psi}(r)|(1+r)\lesssim\frac{1+r}{1+r^{2}}\leq 1.$
@@
$|\hat{\psi}(r)||e^{-2\pi iy\cdot\frac{\xi}{|\xi|}}-2\pi iy\cdot\frac{\xi}{|\xi|}-1|\lesssim|y|^{2}\leq r^{2}.$
@@
$\lesssim\int_{0}^{1}r^{d+4}r^{-d-1-2\alpha}\,\mathrm{d}r+\int_{1}^{\infty}\frac{\,\mathrm{d}r}{r^{1+2\alpha}}<\infty.$
@@
(3.13) $\int_{0}^{R}\int_{B_{R}(z)}\left(\frac{\overline{\nu}_{\lfloor\alpha\rfloor}^{f}(x,r)}{r^{\alpha-1}}\right)^{2}\frac{\,\mathrm{d}x\,\mathrm{d}r}{r}\leq CR^{d},$
@@
$g(x):=\phi\bigg{(}\frac{x-z}{R}\bigg{)}\big{(}f(x)-f(z)+\nabla(f\ast\varphi_{R})(z)\cdot(x-z)\big{)}.$
@@
$\int_{0}^{R}\int_{B_{R}(z)}\left(\overline{\nu}_{1}^{f}(x,r)\right)^{2}\frac{\,\mathrm{d}x\,\mathrm{d}r}{r}=\int_{0}^{R}\int_{B_{R}(z)}\left(\overline{\nu}_{1}^{g}(x,r)\right)^{2}\frac{\,\mathrm{d}x\,\mathrm{d}r}{r}.$